# The efficiency of black hole formation via collisions in stellar systems:

### v1

@@ Appendix B Cluster mass loss
The relation between the initial mass $M_{\mathrm{i}}$ and final mass $M_{\mathrm{f}}$ of stellar systems is not always well-known. Simulations studying the formation of massive objects via collisions usually report the initial stellar mass of the system as well as the mass of the massive object that forms; sometimes also additional information on the final stellar mass and the escapers is being reported. Observations on the other hand only report about the current state of a stellar system. For the analysis in this paper, the relation between final and initial stellar mass can however be somewhat relevant and is therefore examined here in some detail.
As a first and very simple approximation, we could assume that the total mass in the system is conserved, i.e. there are no escapers, and the mass exchange happens only between the stellar mass and the central massive object as a result of collisions. In this case, we have
@@
In the case of simulations, $M_{\mathrm{i}}$ and $M_{\mathrm{BH}}$ are known. The critical mass $M_{\mathrm{crit}}$ can be calculated from the initial conditions and $M_{\mathrm{f}}$ can be estimated from Eq. 5. Employing this approximation, this allows us to plot $M_{\mathrm{f}}/M_{\mathrm{i}}$ as a function of $M_{\mathrm{i}}/M_{\mathrm{crit}}$ for the simulations, which we provide in the top panel of Fig. 4. Under this assumption, we see clearly that for small ratios of $M_{\mathrm{i}}/M_{\mathrm{crit}}\lesssim 0.1$, $M_{\mathrm{f}}/M_{\mathrm{i}}\sim 1$, while it decreases for higher values of $M_{\mathrm{i}}/M_{\mathrm{crit}}$. This behaviour should indeed be expected if the mass of the most massive object increases with $M_{\mathrm{i}}/M_{\mathrm{crit}}$.
We can test the approximation employed in Eq. 5 at least for some simulations where both the initial and final stellar masses are available. For this purpose, we employ the numerical simulations by Vergara et al. (2021, 2023) introduced above, for which we can plot the relation both using the correct values from the simulation, and comparing with the result when instead using the approximation from Eq. 5. In the bottom panel of Fig. 4, we find that the results in both cases are qualitatively similar, though when using the real data, the ratio $M_{\mathrm{f}}/M_{\mathrm{i}}$ decreases somewhat more rapidly with increasing $M_{\mathrm{i}}/M_{\mathrm{crit}}$ due to the escapers. This result also intuitively makes sense, as a higher ratio of $M_{\mathrm{i}}/M_{\mathrm{crit}}$ implies a higher probability that interactions within the cluster have occurred, which can also lead to the ejection of stars from the system. From the data obtained here, we derive a polynomial fit to determine $M_{\mathrm{f}}/M_{\mathrm{i}}$ as a function of $M_{\mathrm{i}}/M_{\mathrm{crit}}$. Our fit in the first case when employing Eq. 5 is given as
@@
Figure 4: Top: $M_{\mathrm{f}}/M_{\mathrm{i}}$ as a function of $M_{\mathrm{i}}/M_{\mathrm{crit}}$, assuming that $M_{\mathrm{i}}=M_{\mathrm{f}}+M_{\mathrm{BH}}$ for all simulations considered from Table 1. Bottom: $M_{\mathrm{f}}/M_{\mathrm{i}}$ as a function of $M_{\mathrm{i}}/M_{\mathrm{crit}}$ using the simulations of Reinoso et al. (2020); Vergara et al. (2021, 2023). We show both the result as it would appear considering the approximation used in Eq. 5 via the solid line, as well as the result when we employ the initial and final stellar masses as obtained from the simulation. In the second case the plot is qualitatively similar but with a steeper decrease due to the escapers. The figure also indicates our polynomial fits to the obtained relations.
From the ratio between these two expressions, we can further derive a scale factor
$\alpha=\left(M_{\mathrm{f}}/M_{\mathrm{i}}\right)_{\rm noesc}-\left(M_{\mathrm{f}}/M_{\mathrm{i}}\right)_{\rm esc}$ (8)
that we can be employed as an approximate correction to estimate the effect of escapers when the information is not given. This procedure allows us to precisely take into account the mass exchange between mass in stars and the mass of the most massive object according to the simulation results, while the mass loss due to escapers is only estimated. The result of this estimate is given in Fig. 5. We note that it shows a similar behaviour as found in Fig. 4, which in principle is natural as the basic trend has been there even when not considering mass loss due to escapers, which is then only further enhanced when that possibility is taken into account.

### v2

@@ Appendix B Cluster mass loss
The relation between the initial mass, $M_{\mathrm{i}}$, and final mass, $M_{\mathrm{f}}$, of stellar systems is not always well known. Simulations studying the formation of massive objects via collisions usually report the initial stellar mass of the system as well as the mass of the massive object that forms; sometimes, also additional information on the final stellar mass and the escapers is being reported. Observations, on the other hand, only report on the current state of a stellar system. For the analysis in this paper, the relation between final and initial stellar mass can, however, be somewhat relevant and, therefore, this aspect is examined here in some detail.
As a first and very simple approximation, we could assume that the total mass in the system is conserved, namely, that there are no escapers and the mass exchange happens only between the stellar mass and the central massive object as a result of collisions. In this case, we have:
@@
In the case of the simulations, $M_{\mathrm{i}}$ and $M_{\mathrm{BH}}$ are known. The critical mass, $M_{\mathrm{crit}}$, can be calculated from the initial conditions and $M_{\mathrm{f}}$ can be estimated from Eq. 5. Employing this approximation, this allows us to plot $M_{\mathrm{f}}/M_{\mathrm{i}}$ as a function of $M_{\mathrm{i}}/M_{\mathrm{crit}}$ for the simulations, which we provide in the top panel of Fig. 4. Under this assumption, we see clearly that for small ratios of $M_{\mathrm{i}}/M_{\mathrm{crit}}\lesssim 0.1$, $M_{\mathrm{f}}/M_{\mathrm{i}}\sim 1$, while it decreases for higher values of $M_{\mathrm{i}}/M_{\mathrm{crit}}$. This behavior should indeed be expected if the mass of the most massive object increases with $M_{\mathrm{i}}/M_{\mathrm{crit}}$.
We can test the approximation employed in Eq. 5, at least for some simulations where both the initial and final stellar masses are available. For this purpose, we can employ the numerical simulations by Vergara et al. (2021, 2023) introduced above, for which we plot the relation both using the correct values from the simulation, and comparing with the result when instead using the approximation from Eq. 5. In the bottom panel of Fig. 4, we show that the results in both cases are qualitatively similar, however, when using the real data, the ratio $M_{\mathrm{f}}/M_{\mathrm{i}}$ decreases somewhat more rapidly with increasing $M_{\mathrm{i}}/M_{\mathrm{crit}}$ due to the escapers. This result also intuitively makes sense, as a higher ratio of $M_{\mathrm{i}}/M_{\mathrm{crit}}$ implies a higher probability that interactions within the cluster have occurred, which can also lead to the ejection of stars from the system. From the data obtained here, we derive a polynomial fit to determine $M_{\mathrm{f}}/M_{\mathrm{i}}$ as a function of $M_{\mathrm{i}}/M_{\mathrm{crit}}$. Our fit in the first case when employing Eq. 5 is given as
@@
Figure 4: $M_{\mathrm{f}}/M_{\mathrm{i}}$ as a function of $M_{\mathrm{i}}/M_{\mathrm{crit}}$, assuming that $M_{\mathrm{i}}=M_{\mathrm{f}}+M_{\mathrm{BH}}$ for all simulations considered from Table 1 (top). $M_{\mathrm{f}}/M_{\mathrm{i}}$ as a function of $M_{\mathrm{i}}/M_{\mathrm{crit}}$ (bottom) using the simulations of Reinoso et al. (2020); Vergara et al. (2021, 2023). We show both the result as it would appear considering the approximation used in Eq. 5 via the solid line and the result when we employ the initial and final stellar masses obtained from the simulation. In the second case the plot is qualitatively similar but with a steeper decrease due to the escapers. The figure also indicates our polynomial fits to the obtained relations.
From the ratio between these two expressions, we can further derive a scale factor of
$\alpha=\left(M_{\mathrm{f}}/M_{\mathrm{i}}\right)_{\rm noesc}-\left(M_{\mathrm{f}}/M_{\mathrm{i}}\right)_{\rm esc},$ (8)
which can be employed as an approximate correction to estimate the effect of escapers when the information is not given. This procedure allows us to precisely take into account the mass exchange between mass in stars and the mass of the most massive object according to the simulation results, while the mass loss due to escapers is only estimated. The result of this estimate is given in Fig. 5. We note that it shows a similar behavior as found in Fig. 4, which, in principle, is natural as the basic trend is seen even when the mass loss due to escapers is not considered; this trend is then further enhanced when that possibility is taken into account.